On January 1, Tolson Company purchased a building by paying $85,000. The building has an estimated life of 40 years and an estimated residual value of $5,000. At December 31, for what gross value will the building appear on the balance sheet? How much accumulated depreciation will be reported at that same date? What will be the net book value of the building on December 31?
Answer #1

First Journal 

Jan 1 Building  - Debit $85,000    

Cash - Credit $85,000


Second Journal 

Dec 31 Depreciation expense  Debit $2000

Accumulated Depreciation Credit $2000


Straight-line depreciation - Subtract the salvage value from the asset cost to get total depreciation, then divide that by useful life to get annual depreciation. 

85,000 - 5,000= 80,000/12=2,000

source: Financial accounting
Answer #2

Gross Value at Dec 31 = $ 85,000

Depreciation per annum = (85,000 - 5,000)/40 = $ 2,000

Accumulated Depreciation at Dec 31 = $ 2,000

Net Value on Dec 31 = Gross Value - Accumulated Depreciation = 85,000 - 2,000

= $ 83,000